Solve the equation by using the square root property. Express any complex numbers using notation.

(2x+5)^2=49

Answer by RAY100(1637)

(2x +5)^2 = 49
.
square root of both sides
.
(2x+5) = +/- 7
.
(1) 2x+5 = +7
.
2x = 2
.
x=1
.
(2) 2x +5 = -7
.
2x = -12
.
x=-6
.
checking
x=1,,, (2(1) +5)^2 = 7^2 =49,,,,ok
.
x= -6,,,(2(-6) +5)^2 = (-12 +5)^2 = (-7)^2 = 49,,,,ok
.
Remember problems to 2nd power have 2 answers